# A Comprehensive Survey on Knowledge Distillation

### v1

@@ VI-E Visual Recognition Distillation
Knowledge distillation has been widely applied to computer vision tasks, enabling the training of compact student models replicating the behavior of larger, resource-intensive teacher models. Key applications of KD in visual recognition are summarized in Table XI, based on two main factors: the type of knowledge transferred (feature-based, logit-based, similarity-based, and combinations) and the distillation scheme employed (offline, online, and self-distillation). Several important visual recognition tasks are covered in this table, including depth estimation [656, 657, 658, 659, 660, 661, 662, 663, 664, 665, 666, 667, 668, 669, 670, 671], face recognition [672, 673, 674, 675, 676, 677, 678, 679, 680, 681, 682, 683], image segmentation [684, 70, 685, 595, 77, 76, 67, 686, 196, 687, 688, 689, 59, 57, 690, 251, 691, 692, 693, 694, 54, 40, 695], medical image analysis [696, 697, 698, 699, 700, 701, 702, 703, 704, 705, 706, 707, 708, 709, 710, 711, 265, 712, 713, 714, 715, 716], object detection [717, 718, 719, 720, 721, 722, 70, 723, 724, 507, 725, 726, 727, 686, 728, 729, 730, 731, 732, 733, 734, 735, 736, 737, 738, 739, 740, 741, 742, 743, 744], object tracking [745, 746, 747, 748], pose estimation [749, 750, 751, 752, 753, 754, 755, 756, 93, 757, 758, 759, 760], super resolution [761, 262, 762, 763, 764, 765, 766, 767, 768, 769, 770, 771, 772, 773], action recognition [774, 431, 214, 775, 776, 777, 778, 779, 780], and image retrieval [781, 782, 783, 784, 785, 786, 787, 788, 789, 790].
@@
TABLE XI: Summary of knowledge distillation methods in visual recognition tasks based on their sources and schemes of distillation.
Task Knowledge Distillation Scheme Reference
Depth Estimation feature offline [656]
self-distillation [657]
logit offline [658], [659], [660], [661], [662]
self-distillation [663], [664], [665], [666], [667], [668]
similarity offline [669]
feature + logit offline [670]
online [671]
Face Recognition feature offline [672], [667], [673], [674], [675], [676], [677], [678]
logit offline [679], [680], [681]
similarity offline [682]
self-distillation [683]
@@
Image Segmentation feature offline [684], [70], [685], [595], [77], [76], [67]
self-distillation [686]
logit offline [196], [659], [687]
online [688], [689]
similarity offline [59], [57]
online [690]
feature + logit offline [251], [691], [692], [693]
online [694]
feature + similarity offline [54], [40]
online [695]
Medical Image Analysis feature offline [696], [697], [698], [699], [700], [701], [702], [703]
online [704]
self-distillation [705], [706]
logit offline [707], [708]
online [709]
self-distillation [710], [711]
similarity online [265]
self-distillation [712], [713]
feature + logit offline [714]
feature + similarity offline [715], [716]
@@
Object Detection feature offline [717], [718], [719], [720], [721], [722], [70], [723], [724], [507], [725]
self-distillation [726], [727], [686], [728]
logit offline [729], [730], [731]
self-distillation [732]
online [733]
similarity offline [734]
feature + logit offline [735], [736]
online [737], [738], [739], [740], [741], [742]
feature + similarity offline [737], [738], [739], [740], [741], [742]
self-distillation [743]
logit + similarity offline [744]
Object Tracking feature offline [745], [746]
logit offline [747]
feature + similarity offline [748]
Pose Estimation feature offline [749]
self-distillation [750]
logit offline [751], [752], [753], [754], [755], [756]
online [93]
self-distillation [757]
feature + logit offline [758], [759]
self-distillation [760]
Super Resolution feature offline [761], [262]
self-distillation [762], [763]
logit offline [764]
self-distillation [765]
relation offline [766], [767]
feature + similarity offline [768], [769]
self-distillation [770]
feature + logit offline [771], [772], [773]
@@
Action Recognition feature offline [774], [431], [214], [775], [776]
online [777]
logit offline [778]
feature + similarity offline [779], [780]
Image Retrieval feature offline [781], [782], [783]
similarity offline [784], [785], [786], [787]
feature + logit offline [788]
feature + similarity offline [789], [790]
VII Performance Comparison
@@
TABLE XII: Performance comparison of different knowledge distillation methods on CIFAR-100 and CityScapes datasets (* denotes that results are not reported from the original paper).
Classification (CIFAR-100)

### v2

@@ 6.5 Visual Recognition Distillation
Knowledge distillation has been widely applied to computer vision tasks, enabling the training of compact student models replicating the behavior of larger, resource-intensive teacher models. Key applications of KD in visual recognition are summarized in Table LABEL:tab:paper-classification-reordered, based on two main factors: the type of knowledge transferred (feature-based, logit-based, similarity-based, and combinations) and the distillation scheme employed (offline, online, and self-distillation). Several important visual recognition tasks are covered in this table, including depth estimation, face recognition, image segmentation, medical imaging, object detection, object tracking, pose estimation, super resolution, action recognition, and image retrieval.
@@
Table 11: Summary of knowledge distillation methods in visual recognition tasks based on their sources and schemes of distillation.
Task Knowledge Scheme Reference
Depth Estimation feature offline (López-Cifuentes et al., 2023)
self-distillation (Zhou & Dong, 2022)
logit offline (Hu et al., 2023b), (Xu et al., 2023c), (Shi et al., 2023), (Ka et al., 2024), (Wang & Liu, 2024)
self-distillation (Peng et al., 2021), (Zhao et al., 2023b), (Marsal et al., 2024), (Hu et al., 2024a), (Boutros et al., 2024), (Dong et al., 2024c)
similarity offline (Chen et al., 2024e)
feature + logit offline (Wu et al., 2023f)
online (Shao et al., 2024b)
Face Recognition feature offline (Shin et al., 2022b), (Boutros et al., 2024), (Otroshi-Shahreza et al., 2023), (Neto et al., 2024), (Caldeira et al., 2024), (Li et al., 2023b), (Huang et al., 2024f), (Babnik et al., 2024)
logit offline (Jung et al., 2021), (Liu et al., 2021a), (Xu et al., 2023a)
similarity offline (Huang et al., 2022d)
self-distillation (Di et al., 2024)
@@
Image Segmentation feature offline (Ji et al., 2022), (Yang et al., 2022f), (Kang et al., 2023a), (Shang et al., 2023a), (Yuan et al., 2024a), (Liu et al., 2024g), (Shu et al., 2021)
self-distillation (Wu et al., 2024b)
logit offline (Amirkhani et al., 2021), (Xu et al., 2023c), (Li et al., 2024c)
online (Shen et al., 2023), (Berrada et al., 2024)
similarity offline (Yang et al., 2022c), (Mansourian et al., 2025)
online (Gao et al., 2022a)
feature + logit offline (Tian et al., 2022c), (Phan et al., 2022), (Xiao et al., 2023a), (Zheng et al., 2025)
online (Zhu et al., 2023a)
feature + similarity offline (Feng et al., 2021), (Liu et al., 2024d)
online (Wang et al., 2024p)
Medical Imaging feature offline (Qin et al., 2021), (Li et al., 2022d), (Ghosh et al., 2023), (Wang et al., 2023g), (Cao et al., 2024), (Nasser et al., 2024), (Chen et al., 2024d), (Zhou et al., 2024c)
online (Dong et al., 2025)
self-distillation (Ye et al., 2022), (Yi et al., 2024a)
logit offline (Feng et al., 2024c), (Wang et al., 2024l)
online (Xie et al., 2023c)
self-distillation (Zhong et al., 2023), (Huang et al., 2024b)
similarity online (Xing et al., 2021)
self-distillation (You et al., 2022), (Sharma et al., 2024)
feature + logit offline (El-Assiouti et al., 2024)
feature + similarity offline (Liu et al., 2022c), (Qi et al., 2024)
@@
Object Detection feature offline (Dai et al., 2021b), (Guo et al., 2021a), (Guo et al., 2021b), (Kang et al., 2021), (Yang et al., 2022a), (Jang et al., 2022), (Yang et al., 2022f), (Lao et al., 2023), (Lan & Tian, 2024), (Wang et al., 2023i), (Liu et al., 2024i)
self-distillation (Wu et al., 2023a), (Jia et al., 2024), (Wu et al., 2024b), (Zheng et al., 2023)
logit offline (Yang et al., 2023e), (Jin et al., 2023a), (Li et al., 2023c)
self-distillation (Zhou et al., 2023a)
online (Nguyen et al., 2022)
similarity offline (Ni et al., 2023)
feature + logit offline (Tang et al., 2022a), (Du et al., 2021)
online (Zhang & Ma, 2021), (Yao et al., 2021), (Yang et al., 2022e), (Li et al., 2022b), (Zhang & Ma, 2023), (Zhu et al., 2023c)
feature + similarity offline (Zhang & Ma, 2021), (Yao et al., 2021), (Yang et al., 2022e), (Li et al., 2022b), (Zhang & Ma, 2023), (Zhu et al., 2023c)
self-distillation (Zhang et al., 2022c)
logit + similarity offline (Wang et al., 2024f)
Object Tracking feature offline (Hou et al., 2024), (Zhang et al., 2024h)
logit offline (Valverde et al., 2021)
feature + similarity offline (Wang et al., 2024k)
Pose Estimation feature offline (Nie et al., 2019)
self-distillation (Chen et al., 2024f)
logit offline (Guo et al., 2023b), (Ye et al., 2023b), (Wang et al., 2021c), (Aghli & Ribeiro, 2021), (Xu et al., 2022c), (Zhang et al., 2019a)
online (Li et al., 2021b)
self-distillation (Li & Lee, 2021)
feature + logit offline (Fang et al., 2023b), (GUAN et al., 2023)
self-distillation (Yang et al., 2023h)
Super Resolution feature offline (Jiang et al., 2024d), (Liu et al., 2024b)
self-distillation (Wang et al., 2024e), (Wang & Zhang, 2024)
logit offline (Noroozi et al., 2024)
self-distillation (Park, 2024)
relation offline (Angarano et al., 2022), (Zhang et al., 2024j)
feature + similarity offline (Xie et al., 2023a), (Fang et al., 2023a)
self-distillation (Wang et al., 2021f)
feature + logit offline (Zhang et al., 2021g), (Yu et al., 2023), (Wang et al., 2024n)
@@
Action Recognition feature offline (Zhao et al., 2022d), (Wang et al., 2024c), (Lee et al., 2023b), (Huang et al., 2024d), (Hong et al., 2021)
online (Guo et al., 2023a)
logit offline (Radevski et al., 2023)
feature + similarity offline (Dai et al., 2021a), (Bian et al., 2021)
Image Retrieval feature offline (Jin et al., 2020), (Porrello et al., 2020), (Budnik & Avrithis, 2021)
similarity offline (Wu et al., 2022), (Xu et al., 2024a), (Tian et al., 2021), (Xie et al., 2024d)
feature + logit offline (Xie et al., 2023b)
feature + similarity offline (Suma & Tolias, 2023), (Xie et al., 2024c)
7 Performance Comparison
@@
Table 12: Performance comparison of different knowledge distillation methods on the CIFAR-100 classification dataset (* denotes that results are not reported from the original paper).